OpenStudy (anonymous):

For questions 1-2, convert to logarithmic equations. For example, the logarithmic form of "2^3 = 8" is "log ^2 8 = 3". 1. 9 ^(3/2) = 27 2. e^x = 7 Show your work, please. :)

OpenStudy (anonymous):

make the replacements in \(b^x=y\iff \log_b(y)=x\)

OpenStudy (anonymous):

\[b ^{x}\] = y then 9\[^{3/2}\] = 29 = log\[_{9}\](29) = 3/2?

OpenStudy (anonymous):

\(9^{\frac{3}{2}}=27\) here \(b=9, x =\frac{3}{2}, y =27\) so logarithmic form is \[\log_9(27)=\frac{3}{2}\]

OpenStudy (anonymous):

That's what I tried to type but I'm pretty new to this site and can't figure out how to format correctly. I'm sorry. It would be log^9 (27) = 3/2. I don't know why I typed 29. Thanks for your help!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

second one you could write \(\log_e(7)=x\) but instead write \(\ln(7)=x\) because ln means log base e